# Fitts’ List Revisited: An Empirical Study on Function Allocation in a Two-Agent Physical Human-Robot Collaborative Position/Force Task

### v1

@@ Abstract
In this letter, we investigate whether the classical function allocation holds for physical Human-Robot Collaboration, which is important for providing insights for Industry 5.0 to guide how to best augment rather than replace workers. This study empirically tests the applicability of Fitts’ List within physical Human-Robot Collaboration, by conducting a user study (N=26, within-subject design) to evaluate four distinct allocations of position/force control between human and robot in an abstract blending task. We hypothesize that the function in which humans control the position achieves better performance and receives higher user ratings. When allocating position control to the human and force control to the robot, compared to the opposite case, we observed a significant improvement in preventing overblending. This was also perceived better in terms of physical demand and overall system acceptance, while participants experienced greater autonomy, more engagement and less frustration. An interesting insight was that the supervisory role (when the robot controls both position and force control) was rated second best in terms of subjective acceptance. Another surprising insight was that if position control was delegated to the robot, the participants perceived much lower autonomy than when the force control was delegated to the robot. These findings empirically support applying Fitts’ principles to static function allocation for physical collaboration, while also revealing important nuanced user experience trade-offs, particularly regarding perceived autonomy when delegating position control.
KEYWORDS: Physical Human-Robot Interaction, Human Factors and Human-in-the-Loop, Human-Centered Robotics.
1 Introduction
@@
The integration of robots in industrial settings has been growing rapidly in recent years, driven by the need to address complex societal challenges such as labor shortages, aging populations, and ever-increasing production demands. While traditional industrial robots have proven their worth in highly controlled and predictable, high-volume settings like automotive assembly lines, they lack the flexibility required for dynamic environments prevalent in high-mix, low-volume manufacturing or maintenance and repair. As a response, collaborative robots (cobots) have emerged, designed specifically for safe operation near human workers. This unlocks the potential for direct physical Human-Robot Collaboration (pHRC) that promises to merge the complementary skills of both agents, creating more effective, flexible, and meaningful collaborations [1]. Such vision aligns with the human-centric principles of Industry 5.0, which emphasizes technology in service of worker well-being and resilient production systems [2]. However, despite the potential of collaborative robots, achieving truly effective, fluent, and satisfying collaboration remains a significant challenge [3].
@@
One element of this challenge lies in function allocation: deciding which agent–human or robot–should perform which aspects of a shared task [4, 5, 6, 7, 8, 9]. We can model physical interactions with impedance, which deals with two key variables: motion and forces [10]. Consider a task like manual blending or polishing of a complex surface, such as an aircraft fan blade (see Fig. 1a). This requires both precise positioning of the tool and careful application of force. Classic frameworks, notably Fitts’ “MABA-MABA” (Men Are Better At / Machines Are Better At) list stemming from early human factors research [11], provide initial guidance. Despite its age and extensive criticism, this framework has shown remarkable persistence throughout the history of function allocation [12]. The framework suggests humans generally excel at tasks requiring perception, judgment, and adaptability (e.g., identifying and deciding where to polish), while machines excel at speed, consistent force application, and repetitive precision movements (e.g., producing a constant force). Modern cobots offer the technical capability to split control of force and position between the human and robot (see Fig. 1b).
@@
Some existing work in pHRC has exploited this ability of cobots to take human inputs during physical interaction, and allocate positions and force sub-tasks. The approach in [13] made the distribution of functions during an assembly task, where the robot performed most of the large movements, while the human stepped in for precise movements. Humans can also step in with physical interaction to give the robot occasional hints on how to re-plan longer movement trajectories in real-time during assembly, where there are multiple possible goals [14]. The method in [15] was in line with the classical framework, where the assembly of bulky objects was handled by allocating load carrying to the robot, while the precise positioning of the object during installation was allocated to the human. In [16], a collaborative polishing task was solved by allocating force control to the robot and position control to the human, consistent with Fitts’ principle. In industrial tasks like sanding, letting a robot perform repetitive high-force motions can reduce human physical strain, but fully autonomous solutions often struggle with quality and adaptability [17]. Nevertheless, the existing solutions built on the general Fitts’ principle without having a clear insight into whether it is valid in the context of pHRC, prompting a re-evaluation of these classic principles in the context of direct physical collaboration.
@@
To address this gap and provide the missing crucial insight, this study investigates a fundamental question: who should control what in a two-agent physical human-robot collaborative position/force task? To answer this question, we conducted human factors experiments involving a blending task requiring both force application and tool positioning. Participants experienced four distinct conditions representing the extremes of static role allocation on the shared control spectrum: HH (fully manual control), HR (human controls position, robot controls force), RH (robot controls position, human controls force), and RR (robot controls both, human supervises). by collecting and statistically analyzing both objective and subjective data across these conditions, we aimed to gain new insights into function allocation in the context of pHRC. We hypothesize that:
@@
Participants experience lower workload in HR than in RH.
@@
Participants experience greater autonomy, feel more engaged and consider themselves more competent and useful in HR than in RH.
@@ 2.1 Experimental setup
The robot is controlled at the joint-torque level using a Cartesian Impedance Controller. The interaction force/torque in Cartesian space was defined as:
where and are the translational and rotational parts, respectively, of the stiffness matrix . represents the damping matrix, and being the actual and desired robot end-effector pose in Cartesian space, a diagonal matrix used for selecting the axis in which the desired force is applied and the desired force, both in Cartesian space. We used a critically damped design , with damping ratio . The robot and tool were gravity-compensated, while a nullspace controller enforced an ’elbow up’ configuration.
@@
where is the desired end-effector pose at time , is a time-based (spherical) linear interpolation function taking the start pose (), goal pose () and the time divided by the total execution time (), which calculated as follows: , with being the desired velocity.
2.2 Position/force task
@@
2.3 Experimental conditions
The study included four experimental conditions that represent all combinations of extremes on the shared control continuum for both the position and force sub-tasks. The conditions are summarized in Table 1, where also the differences in terms of control with respect to (1) are specified.
@@ 2.3 Experimental conditions
Condition 4 - RR: Robot controls both the position of the tool as well as the application of force, while participants supervise and intervene in case of mistakes made by the robot.
2.4 Experimental design
@@
The four conditions were designed to isolate the effects of delegating position and/or force control. HH provides a fully‑manual baseline, whereas in RR both positioning and force exertion is automated. RR is tuned to operate near the performance limits of what is physically possible for a human to achieve. The two “mixed-role” conditions for robot-assistance include HR, where human controls positioning and the robot the force, and RH, where the human controls the force and the robot the positioning. All conditions used the same Cartesian impedance controller in (1), with condition-specific settings: translational elements of the stiffness matrix were set to N/m in all directions for HH and HR, N/m in and for RH, and N/m in all directions for RR. Rotational stiffness was fixed at Nm/rad in all conditions. Force thresholds were set at N and N. Pixel damage removal time at was s, with health points per pixel. Trajectory velocities (RH/RR) were m/s (blending), m/s ((un)latching), and m/s (travel).
In HR, the robot applied when the tooltip was within 1 mm of the surface, and stopped when it moved above 3 mm, creating an intuitive latching effect. In RR, the robot autonomously controlled both position and force by following a predefined adaptive trajectory that systematically covered the damage spot, divided into sections of equal width. If insufficient removal was detected in a section, the robot would revisit it. To make the task engaging for the human in the supervisory role, at fixed points, autonomy errors were introduced by deviating the trajectory outside the damaged area, leading to overblending unless the participant intervened by pressing a button. This condition modeled supervisory control, acknowledging that full autonomy is infeasible in most complex practical tasks [18].
2.5 Procedure
@@
2.6 Quantitative measures
The objective performance metrics that were considered are the time of completion, which is defined as the amount of time it takes the participant to remove the damage during a trial, and overblended health defined as the amount of cumulative health points resulting from applying force at locations where there is no damage (any longer). Participants’ perceived acceptance was assessed using the Van der Laan usefulness and satisfying scales [19], while perceived workload was assessed using the NASA-TLX questionnaire [20].
Recognizing the need to also probe the psychological impact of pHRC beyond standard metrics [21], we developed the Agentic Collaboration and Engagement Scale (ACES). This scale targets four constructs–autonomy, competence, usefulness, and engagement–identified in work-psychology and self-determination theory as drivers of motivation, job satisfaction, and well-being [22]. Assessing these factors is crucial because collaborative robots can reshape meaningful work [23] and key aspects of job quality [24]. ACES employs a dual-perspective format rating each construct twice: once for the participant’s own experience “I”) and once for the participant’s perception of the robot (“robot”). The scale consists of the following statements rated on a Likert scale:
@@ 2.6 Quantitative measures
where XX can be “I” or “robot”, resulting in 8 items.
2.7 Statistical evaluation
The statistical evaluation procedure is shown in Fig. 2 G). Data normality was assessed using the Shapiro–Wilk test. If normality was met (, including minor violations), a Repeated Measures ANOVA was used. Violations of sphericity (Mauchly’s test) were corrected with Greenhouse–Geisser. Significant main effects were followed by paired -tests with Holm–Bonferroni correction. Effect size was reported as . If normality was violated (), a Friedman test was applied. Post-hoc Wilcoxon signed-rank tests with Holm–Bonferroni correction were used for pairwise comparisons, and effect size was reported using Kendall’s W.
@@ 2.7 Statistical evaluation
| Completion time | Mdn | 109.33 | 109.99 | 90.72 | 112.87 | (3)=49.71 | W=0.637 | 1: p=0.0528 | 2: p=0.4992 | 3: p0.0001 |
| (s) | IQR | 103.12-111.07 | 103.88-115.77 | 90.09-92.01 | 105.70-119.33 | p0.0001∗∗ | 4: p=0.1261 | 5: p0.0001 | 6: p0.0001 | |
| Overblended health | Mdn | 12368 | 0 | 526921 | 21 | (3)=16.34 | W=0.209 | 1: p=1.0000 | 2: p=0.0705 | 3: p=1.0000 |
| (health points) | IQR | 3467-36513 | 0-11866 | 3993-8156 | 0-9147 | p=0.0010∗∗ | 4: p=0.0397 | 5: p=1.0000 | 6: p0.0001 | |
@@
3 Results
In this section, the experimental results are presented, highlighting the differences across the four conditions (HH, HR, RH, RR). Findings are structured by measure type, referencing Figures 3-6 and the detailed statistics evaluation in Table 2.
3.1 Van der Laan System Acceptance
Significant differences were found across conditions for perceived usefulness () and satisfaction (). Pairwise comparisons revealed a strong preference for HR over RH, with HR rated significantly higher on both usefulness () and satisfaction (). HR was also rated significantly more useful () and satisfying () than the baseline HH condition. Conversely, RH showed no usefulness benefit over HH () and was significantly less satisfying (). Notably, the supervisory RR condition was rated significantly higher than RH on usefulness () and satisfaction (), positioning it second overall in user acceptance, statistically comparable to HH but generally lower than HR.
3.2 NASA-TLX Workload
Significant effects were found for all workload dimensions except mental demand () (Fig. 4). Physical demand showed the largest differences (), with RR being significantly lowest ( vs all). HR significantly reduced physical demand compared to both HH () and RH (). Effort followed similar patterns, being lowest in RR. Frustration was uniquely high in the RH condition, significantly higher than HH (), HR (), and RR (). Although mental demand showed no significant main effect, variability within conditions was high (see SDs in Table 2).
3.3 Performance Metrics
Both completion time () and overblended health () showed significant differences. RR yielded significantly faster completion times than all other conditions (), while cumulative overblended health was significantly higher in RH compared to HR () and RR ().
3.4 Agentic Collaboration and Engagement Scale (ACES)
Significant differences were observed for participants’ ratings of their own autonomy, competence, usefulness, and engagement (all but competence ). Ratings for autonomy, usefulness, and engagement were highest when participants controlled position (HH and HR). Specifically, HR yielded significantly higher ratings compared HR on autonomy (), usefulness (), and engagement (). Delegating position reduced perceived autonomy significantly more (HH vs RH, ) than delegating force (HH vs HR, ). Self-rated competence was lowest in RH, significantly lower than in RR ().
Significant differences were also observed for participants’ ratings of the robot’s autonomy, usefulness, and engagement (all ), but not for competence (). Robot autonomy was perceived highest when it controlled position (RH and RR). The robot in the RR condition was rated significantly most useful and engaging compared to all other conditions.
4 Discussion

### v2

@@ Abstract
In this letter, we investigate whether classical function allocation—the principle of assigning tasks to either a human or a machine—holds for physical Human-Robot Collaboration, which is important for providing insights for Industry 5.0 to guide how to best augment rather than replace workers. This study empirically tests the applicability of Fitts’ List within physical Human-Robot Collaboration, by conducting a user study (N=26, within-subject design) to evaluate four distinct allocations of position/force control between human and robot in an abstract blending task. We hypothesize that the function in which humans control the position achieves better performance and receives higher user ratings. When allocating position control to the human and force control to the robot, compared to the opposite case, we observed a significant improvement in preventing overblending. This was also perceived better in terms of physical demand and overall system acceptance, while participants experienced greater autonomy, more engagement and less frustration. An interesting insight was that the supervisory role (when the robot controls both position and force) was rated second best in terms of subjective acceptance. Another surprising insight was that if position control was delegated to the robot, the participants perceived much lower autonomy than when the force control was delegated to the robot. These findings empirically support applying Fitts’ principles to static function allocation for physical collaboration, while also revealing important nuanced user experience trade-offs, particularly regarding perceived autonomy when delegating position control.
I Introduction
@@
The integration of robots in industrial settings has been growing rapidly in recent years, driven by the need to address complex societal challenges such as labor shortages, aging populations, and ever-increasing production demands. While traditional industrial robots have proven their worth in highly controlled and predictable, high-volume settings like automotive assembly lines, they lack the flexibility required for dynamic environments prevalent in high-mix, low-volume manufacturing or maintenance and repair. As a response, collaborative robots (cobots) have emerged, designed specifically for safe operation near human workers. This unlocks the potential for direct physical Human-Robot Collaboration (pHRC) that promises to merge the complementary skills of both agents, creating more effective, flexible, and meaningful collaborations [ajoudani_progress_2018]. Such vision aligns with the human-centric principles of Industry 5.0, which emphasizes technology in service of worker well-being and resilient production systems. However, despite the potential of collaborative robots, achieving truly effective, fluent, and satisfying collaboration remains a significant challenge [michaelis_collaborative_2020].
@@
One element of this challenge lies in function allocation: deciding which agent–human or robot–should perform which aspects of a shared task [mortl_role_2012, jarrasse_slaves_2014, el_makrini_hierarchical_2022, selvaggio_autonomy_2021, merlo_ergonomic_2023, vianello_effects_2024]. We can model physical interactions with impedance, which deals with two key variables: motion and forces [hogan_impedance_1984]. Consider a task like manual blending or polishing of a complex surface, such as an aircraft fan blade (see Fig. 1 left). This requires both precise positioning of the tool and careful application of force. Classic frameworks, notably Fitts’ “MABA-MABA” (Men Are Better At / Machines Are Better At) list stemming from early human factors research [fitts_human_1951], provide initial guidance. Despite its age and extensive criticism, this framework has shown remarkable persistence throughout the history of function allocation [de_winter_why_2014]. The framework suggests humans generally excel at tasks requiring perception, judgment, and adaptability (e.g., identifying and deciding where to polish), while machines excel at speed, consistent force application, and repetitive precision movements (e.g., producing a constant force). Modern cobots offer the technical capability to split control of force and position between the human and robot (see Fig. 1 right).
@@
Some existing work in pHRC has exploited this ability of cobots to take human inputs during physical interaction, and allocate positions and force sub-tasks. The approach in [cherubini_collaborative_2016] made the distribution of functions during an assembly task, where the robot performed most of the large movements, while the human stepped in for precise movements. Humans can also step in with physical interaction to give the robot occasional hints on how to re-plan longer movement trajectories in real-time during assembly, where there are multiple possible goals [haninger_model_2023]. The method in [roveda_human-robot_2018] was in line with the classical framework, where the assembly of bulky objects was handled by allocating load carrying to the robot, while the precise positioning of the object during installation was allocated to the human. In [peternel_robot_2018], a collaborative polishing task was solved by allocating force control to the robot and position control to the human, consistent with Fitts’ principle. In industrial tasks like sanding, letting a robot perform repetitive high-force motions can reduce human physical strain, but fully autonomous solutions often struggle with quality and adaptability [konstant_humanrobot_2025]. Nevertheless, the existing solutions built on the general Fitts’ principle without having a clear insight into whether it is valid in the context of pHRC, prompting a re-evaluation of these classic principles in the context of direct physical collaboration.
@@
To address this gap and provide the missing insight, this study investigates a fundamental question: who should control what in a two-agent physical human-robot collaborative position/force task? To answer this question, we conducted human factors experiments involving a blending task requiring both force application and tool positioning. Participants experienced four distinct conditions representing the extremes of static task allocation on the shared control spectrum: HH (fully manual control), HR (human controls position, robot controls force), RH (robot controls position, human controls force), and RR (robot controls both, human supervises). By collecting and statistically analyzing both objective and subjective data across these conditions, we aimed to gain new insights into function allocation in the context of pHRC. We hypothesize that:
@@
Perceived workload is lower in HR than in RH.
@@
Perceived autonomy, engagement, competence, and usefulness score higher in HR than in RH.
@@ II-A Experimental Setup
The robot is controlled at the joint-torque level using a Cartesian Impedance Controller operating at 200 Hz. The interaction force/torque in Cartesian space was defined as:
where and are the translational and rotational parts, respectively, of the stiffness matrix . represents the damping matrix, and being the actual and desired robot end-effector pose in Cartesian space, a diagonal matrix used for selecting the axis in which the desired force is applied and the desired force, both in Cartesian space. We used a critically damped design , with damping ratio . The robot and tool were gravity-compensated, while a nullspace controller enforced an “elbow up” configuration.
@@
where is the desired end-effector pose at time , is a time-based interpolation function (linear for position and spherical for orientation) taking the start pose (), goal pose () and the time divided by the total execution time (), which calculated as follows: , with being the desired velocity.
II-B Position/Force Task
@@
II-C Experimental Conditions
The study included four experimental conditions that represent all combinations of extremes on the shared control spectrum for both the position and force sub-tasks. The conditions are summarized in Table I, where also the differences in terms of control with respect to (1) are specified.
@@ II-C Experimental Conditions
Condition 4 - RR: Robot controls both the position of the tool as well as the application of force, while participants supervise and intervene in case the robot makes mistakes.
II-D Experimental Design
@@
The four conditions were designed to isolate the effects of delegating position and/or force control: a manual baseline (HH), an automated supervisory condition (RR), and two “mixed-task” conditions (HR and RH). RR was tuned to operate near the performance limits of what is physically possible for a human to achieve. All conditions used the same Cartesian impedance controller in (1), with condition-specific translational stiffness settings. For HH and HR, translational elements of the stiffness matrix were set to N/m in all directions, to allow for unhindered human positioning. For RH, it was tuned to N/m in and and for RR in all translational directions. Rotational stiffness was fixed at Nm/rad in all conditions. These values were determined through pilot testing to be high enough to rigidly guide the user’s motion along the robot’s path without causing instability. Force thresholds were set at N and N. Pixel damage removal time at was s, with health points per pixel. Trajectory velocities (RH/RR) were tuned to m/s (blending), m/s ((un)latching), and m/s (travel).
In HR, the robot applied a force capped at whenever the tooltip was within 1 mm of the surface and stopped when it moved above or below 3 mm, creating an intuitive latching effect consistent across participants. In RR, the robot autonomously controlled both position and force by following a predefined adaptive trajectory that systematically covered the damage spot, divided into sections of equal width. If insufficient removal was detected in a section, the robot would revisit it. To make the task engaging for the human in the supervisory role, at fixed points, automation errors were introduced by deviating the trajectory outside the damaged area, leading to overblending unless the participant intervened promptly by pressing a button. This condition modeled supervisory control, acknowledging that full autonomy is infeasible in most complex practical tasks [bradshaw_seven_2013].
II-E Procedure
@@
II-F Quantitative Measures
The objective performance metrics that were considered are the time of completion, which is defined as the amount of time it takes the participant to remove the damage during a trial, and overblended health defined as the amount of cumulative health points resulting from applying force at locations where there is no damage (any longer). Participants’ perceived acceptance was assessed using the Van der Laan usefulness and satisfying scales [van_der_laan_simple_1997], while perceived workload was assessed using the NASA-TLX questionnaire [hart_nasa-task_2006].
To probe the psychological impact of pHRC beyond standard metrics [fletcher_we_2023], we developed the exploratory Agentic Collaboration and Engagement Scale (ACES) because existing scales from workplace psychology are ill-suited for a single-session, abstracted experiment and lack a human-robot role perspective. The scale’s theoretical foundation draws from self-determination theory (SDT), which identifies autonomy and competence as basic psychological needs that foster intrinsic motivation and engagement [gagne_understanding_2022]. These factors, together with a sense of usefulness, are drivers of meaningful work, an experience that is known to enhance job satisfaction and well-being [smids_robots_2020]. To assess reliability, Cronbach’s alpha () was calculated, which showed an internal consistency of , 95% CI [0.63, 0.80] for the “Human” scale (measuring perceived meaningful involvement) and , 95% CI [0.71, 0.85] for the “Robot” scale (measuring perceived quality of the robot teammate). ACES uses a dual-perspective format, rating each construct for the participant’s experience (“I”) and their perception of the robot (“the robot”) on a Likert scale using four statements:
@@ II-F Quantitative Measures
II-G Statistical Evaluation
As shown in Fig. 2 G), data normality was assessed using the Shapiro–Wilk test. If normality was met (, including minor violations), a Repeated Measures ANOVA was used. Violations of sphericity (Mauchly’s test) were corrected with Greenhouse–Geisser. Significant main effects were followed by paired -tests with Holm–Bonferroni correction. Effect size was reported as . If normality was majorly violated (), a Friedman test was applied. Post-hoc Wilcoxon signed-rank tests with Holm–Bonferroni correction were used for pairwise comparisons, with effect size reported using Kendall’s W.
III Results
Here, the experimental results are presented, highlighting the differences across the four conditions (HH, HR, RH, RR). Findings are structured by measure type, referencing Figures 3-6 and the detailed statistical evaluation in Table II.
III-A Van der Laan System Acceptance
Significant differences were found across conditions for perceived usefulness () and satisfaction (). Pairwise comparisons revealed a strong preference for HR over RH, with HR rated significantly higher on both usefulness (t-test, ) and satisfaction (t-test, ). HR was also rated significantly more useful (t-test, ) and satisfying (t-test, ) than the baseline HH condition. Conversely, RH showed no usefulness benefit over HH (t-test, ) and was significantly less satisfying (t-test, ). Notably, the supervisory RR condition was rated significantly higher than RH on usefulness (t-test, ) and satisfaction (t-test, ), positioning it second in user acceptance, statistically comparable to HH but generally lower than HR.
III-B NASA-TLX Workload
Significant effects were found for all workload dimensions except mental demand () (Fig. 4). Physical demand showed the largest differences (), with RR being significantly lowest (t-test, ) compared to all other conditions. HR significantly reduced physical demand compared to both HH (t-test, ) and RH (t-test, ). Effort followed similar patterns, being lowest in RR. Frustration was uniquely high in the RH condition, significantly higher than HH (t-test, ), HR (t-test, ), and RR (t-test, ). Although mental demand showed no significant main effect, variability within conditions was high (see SDs in Table II).
III-C Performance Metrics
Both completion time () and overblended health () showed significant differences. RR yielded significantly faster completion times than all other conditions (Wilcoxon signed-rank, ), while overblended health was significantly higher in RH compared to HR (Wilcoxon signed-rank, ) and RR (Wilcoxon signed-rank, ).
III-D Agentic Collaboration and Engagement Scale (ACES)
Significant differences were observed for participants’ ratings of their own autonomy ((3)=64.78, ), competence ((3)=17.83, , usefulness (), and engagement (). Ratings for autonomy, usefulness, and engagement were highest when participants controlled position (HH and HR). Specifically, HR yielded significantly higher ratings compared HR on autonomy (Wilcoxon signed-rank, ), usefulness (t-test, ), and engagement (t-test, ). Delegating position reduced perceived autonomy significantly more (HH vs RH: Wilcoxon signed-rank, ) than delegating force (HH vs HR: Wilcoxon signed-rank, ). Self-rated competence was lowest in RH, significantly lower than in RR (Wilcoxon signed-rank, ).
Significant differences were also observed for ratings of the robot’s autonomy ((3)=60.75, ), usefulness (), and engagement (), except for competence (). Robot autonomy was perceived highest when it controlled position (RH and RR). The robot in the RR condition was rated significantly most useful and engaging compared to all other conditions.
@@ III-D Agentic Collaboration and Engagement Scale (ACES)
| Completion time | Mdn | 109.99 | 112.87 | 109.33 | 90.72 | (3)=49.71 | W=0.637 | 1: p=0.0528 | 2: p=0.4992 | 3: p0.0001 |
| (s) | IQR | 103.88-115.77 | 105.70-119.33 | 103.12-111.07 | 90.09-92.01 | p0.0001∘∘ | 4: p=0.1261 | 5: p0.0001 | 6: p0.0001 | |
| Overblended health | Mdn | 0 | 21 | 12368 | 5269 | (3)=16.34 | W=0.209 | 1: p=1.0000 | 2: p=0.0705 | 3: p=1.0000 |
| (health points) | IQR | 0-11866 | 0-9147 | 3467-36513 | 3993-8156 | p=0.0010∘∘ | 4: p=0.0397 | 5: p=1.0000 | 6: p0.0001 | |
@@
| : Higher values are better, : lower values are better Color scale (normalized for each metric): | ||||||||||
@@
IV Discussion